Template for 18.904

Mark Behrens November 29, 2005

1 Introduction

I wrote this template to remind you guys how TEXworks. I’ve included some things that I like to use, but you probably have your own preferences. Definition 1.1. This command was defined in the header. Notice that it numbers according to the section Definition. However, you may not want your definition numbered.

Equations, the way I’ve set it up, number combadibly with everything else.

e^{2πi} − 1 = 0  (1.2)

Note that the number works well with Definition 1.1. Also note how I used the label and ref commands to reference the definition in the last sentence. However, you may want equations without numbers.

ax + b = c

And sometimes you want inline equations like ∫_{M} x^{2} dx.

2 Other info

Theorem 2.1. Let A be the set of all sets. Then A is really big. Proof. This is a trivial consequence of general nonsense.

Check out the reference [1].

References

[1] Oetiker, et. al. The not so short introduction to Latex 2e.